OpenStudy (anonymous):

Solve the quadratic equation: 3x2 − 2x − 8 = 0

OpenStudy (anonymous):

3x^2-2x-8 3x^2-6x-4x-8 3x(x-2)-4(x-2)=0 (3x-4)(x-2)=0 x=4/3, x=2